Complete the following short-run cost table using the information provided.

Total

product

TFC

AFC

TVC

AVC

TC

MC

0

0.0

1

3.0

2

5.0

5.0

3

9.0

7. (10 points) Answer the questions below on the basis of the above graph.

(a) (3 points-1 point for short run; 1 point for long run; 1 point for how you can tell) How can you tell if these cost curves are for the short run or the long run?

(b) (7 points) Based on the graph answer the following questions. (1 point each)

(1) AVC at 6,000 units of output?

(2) ATC at 6,000 units of output?

(3) AFC at 6,000 units of output?

(4) TVC at 6,000 units of output?

(5) TFC at all levels of output?

(6) TC at 10,000 units of output?

(7) When diminishing marginal productivity sets in?


you need to fill in the table thats the question
Answer #1

(a)

When Q = 2, AFC = 5 which means total fixed cost (TFC) is positive. Since fixed costs exist only in short run, this is a short run scenario.

(b)

  • TFC = AFC x Q = 5 x 2 = 10 and this remains unchanged across all output levels
  • AFC = TFC/Q
  • AVC = TVC/Q
  • TC = TFC + TVC
  • ATC = TC/Q
  • MC = Change in TC / Change in Q
Total product (Q) TFC AFC TVC AVC TC ATC MC
0 10 0 10
1 10 10 3 3 13 13 3
2 10 5 5 2.5 15 7.5 2
3 10 3.33 9 3 19 6.33 4